# Fundamental physical constants, operation of physical phenomena and entropy increase

### v5

@@ Abstract
1 Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. [1, 2, 3, 4]), and their recommended values are updated and maintained in the National Institute for Standards and Technology database [5].
@@ 1 Introduction
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future [6, 7, 8, 9, 10, 11, 12, 13, 3]. The “barcodes of ultimate reality” [6], FPCs also set our Universe apart from those worlds where these constants might be different.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science [14]. The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental [15]. We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed [6, 7, 8, 9, 10, 11, 12, 13]? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks [16, 12].
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational and cosmological constants. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form [6, 13].
@@
For the rest of intermediate length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life [6, 13]. The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. [13] discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC. FPCs involved in different processes at intermediate scales are often the same and include the Planck constant, speed of light in vacuum, electron mass and charge and the proton-to-electron mass ratio (this is different from FPCs operating at the very top and bottom scales mentioned earlier - these play role at their length and energy scales only). This means that the parameter space of these FPCs can be better constrained and understood.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity [17], diffusion and spin dynamics [18, 19, 20, 21, 22], electron and heat transport [23, 24, 25, 26, 27], thermal conductivity in insulators [28] and conductors [29], the speed of sound [30], elastic moduli including those in low-dimensional systems [29, 31], flow in cells [32, 33], vibration frequency in solids and superconducting temperature [34]. This has deepened our understanding of these properties and their physical models.
@@
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy. We propose that this is a general effect operating at different levels of structure in the Universe.
The entropy increase enabled by FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution. Based on the entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussion [35, 36, 37, 38, 39, 40, 41, 42, 6, 13, 33].
2 Why filling the condensed matter gap is important
@@
Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs [13]. On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@
In the next sections, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
3 Observability and operation of phase transitions. Emergence of new states with higher entropy
3.1 Fundamental and environmental parameters
@@
We will be discussing FPCs central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$ and the proton mass $m_{p}$. As discussed in this section, the first three constants set the characteristic scales of distance and energy in condensed matter. $m_{p}$ sets the scale of lattice oscillation frequency (see section 3.2.2). We will also be discussing two dimensionless ratios, the proton-to-electron mass ratio $\beta$
$\beta=\frac{m_{p}}{m_{e}}$ (1)
and the fine-structure constant
$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (2)
where $c$ is the speed of light in vacuum.
$\alpha$ and $\beta$ are important in key processes at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) [6, 13].
It is useful at this stage to recall that characteristic scales of distance and energy in condensed matter are set by the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
@@
3.2 Structural phase transitions
3.2.1 Entropy increase due to a phase transition
Our first case study are structural phase transitions in solids. Let’s consider common silica SiO2 which has several polymorphs at different pressure and temperature, separated by phase transitions. As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase [44, 45].
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is [46]:
@@ 3.2.1 Entropy increase due to a phase transition
As expected for high-temperature phases on general grounds [46], cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra [44, 45]. This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (5).
@@
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are [45, 47]:
@@
The entropy increase can also take place below $T_{c}$ if the high-temperature phase is a long-lived metastable state below $T_{c}$. In our quartz-cristobalite example, the phase transition is a reconstructive transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ [44, 45]. We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics [46]. This is discussed in Section 3.5 in more detail.
@@
The quartz-cristobalite example is readily generalised to many systems where the combined entropy increases due to phase transitions. Figure 2 illustrates this point.
Figure 2: Quartz on the left transforms to cristobalite at high temperature as a result of a phase transition. In a set of solids, phase transitions result in larger phase space and higher entropy. Shown minerals are for illustration only.
3.2.2 Observability of phase transitions. Entropy changes.
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of fundamental constants [34]. This bound follows from combining Eq. (4) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$ [17]
@@
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as [34]
@@
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (1). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 [6, 13].
Let’s consider what would happen to a phase transition were FPCs to take different values. Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (12)-(13). Note that this question is well-posed because it is possible to vary $\omega_{m}^{u}$ by varying $m_{e}$ and $m_{p}$ or $\hbar$ and $e$ while keeping $\beta$ and $\alpha$ in Eqs. (1) and (2) unchanged and hence maintain all preceding essential processes (e.g. stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact [6, 13].
@@
Recalling that $\omega_{m}$ is close to $\omega_{\rm D}$, condition (14) implies $T\ll\hbar\omega_{\rm D}$. This, in turn, implies that the high-temperature classical result (5) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (5), the entropy is [46]
@@
The observability of phase transitions can also be discussed on the basis of $T_{c}$. Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles [45]. This energy is close to the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (12),
@@
$T_{c}$ can be estimated by recalling that $A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1-15, with an average of 8. Using this value and $\omega_{m}^{u}=3680$ K (13), Eq. (17) gives $T_{c}\approx 500$ K. This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory [48] and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
@@
In summary, the discussion in Section 3.2 has so far given us several insights:
@@
    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections 3.3, 3.4 and 3.5).
@@
    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections 3.4, 3.5 and 4 below).
3.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked [34] whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (13) [34]. $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (13), this gives $T_{c}^{u}$ as [34]

### v6

@@ Abstract
I Introduction
Fundamental physical constants (FPCs) determine physical processes at all energy and length scales of our experiments and theories. These different length scales are illustrated in Figure 1. The values of FPCs are often listed in textbooks and reviews (see, e.g., Refs. Tiesinga et al. (2021); Ashcroft and Mermin (1976); Uzan (2003, 2011)), and their recommended values are updated and maintained in the National Institute for Standards and Technology database NIST .
@@ I Introduction
Together with our fundamental theories, FPCs are the key to understanding the present state of our Universe and its contents as well as its past and future Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019); Uzan (2003). The “barcodes of ultimate reality” Barrow (2003), FPCs also set our Universe apart from those worlds where these constants might be different.
@@
Understanding the values and origin of FPCs is one of the grandest challenges in modern science Barrow and Webb (2006). The problem of understanding FPCs is not that it involves very hard technical calculations but that we do not know anything more fundamental Weinberg (1983). We do not know what kind of theoretical tools we need to address the problem. It is probably this enormity of the problem that explains why there seems to be no sustained programme of research in this area, scheduled meetings and so on, as compared to more developed fields of science.
@@
A starting point to understand FPCs is to ask what would happen were FPCs to take values different from those currently observed Barrow (2003); Barrow and Tipler (2009); Carr and Rees (1979); Carr (2009); Sloan et al. (2022); Cahn (1996); Hogan (2000); Adams (2019)? This is done on the basis of our physical models. The bottom in Figure 1 involves length scales of particle physics. According to our theories, observed particles involve a finely-tuned balance between the masses of down and up quarks: larger up-quark mass gives the neutron world without protons and hence no atoms made from nuclei and electrons around them, whereas larger down-quark mass gives the proton world without neutrons where light atoms can form only but not heavy atoms. Our world with heavy atoms and electrons endowing complex chemistry would not exist with a few per cent fractional change in the mass difference of the two quarks Hogan (2009, 2000).
@@
The top in Figure 1 involves cosmological scales and a balance between the gravitational and cosmological constants. This balance gives enough time for the Universe to expand and not to contract too quickly and at the same time maintain sufficient gravitational clamping so that galaxies, stars and other structures can form Barrow (2003); Adams (2019).
@@
For the rest of intermediate length scales in Figure 1, our theories imply that there is an intricate network of inter-related constraints on FPCs that need to be satisfied to give many structures in the observable Universe. This includes formation and evolution of different astronomical objects and in particular stars, synthesis of heavy elements in stars including carbon and oxygen and other elements which can then form molecular structures essential to life Barrow (2003); Adams (2019). The strength of these constraints, the degree of tuning between FPCs, depends on how well developed our physical models are. Ref. Adams (2019) discusses existing uncertainties involved in our models and how these uncertainties affect constraints imposed on FPC.
@@
It was realised that FPCs set upper or lower bounds on several key condensed matter properties. Examples include bounds on viscosity Trachenko and Brazhkin (2020), diffusion and spin dynamics Luciuk et al. (2017); Sommer et al. (2011); Bardon et al. (2014); Trotzky et al. (2015); Enss and Thywissen (2018), electron and heat transport Hartnoll (2015); Mousatov and Hartnoll (2020); Behnia and Kapitulnik (2019); Nussinov and Chakrabarty (2022); Grozdanov (2021), thermal conductivity in insulators Trachenko et al. (2021) and conductors Brazhkin (2023), the speed of sound Trachenko et al. (2020), elastic moduli including those in low-dimensional systems Brazhkin (2023); Trachenko (2023a), flow in cells Trachenko (2023b, c), vibration frequency in solids and superconducting temperature Trachenko et al. . This has deepened our understanding of these properties and their physical models.
@@
We show that the new states emerging from various transitions increase the phase space and entropy. We also show that were FPCs to take different values, these transitions would become inoperative at our environmental conditions, and the new states would not emerge. This suggests that the current values of FPCs, by enabling various transitions and reactions which produce the new states, result in the increase of the phase space and entropy. We propose that this is a general effect operating at different levels of structure in the Universe. The entropy increase due to FPCs can be discussed in terms of an evolution-like mechanism where this increase is analogous to selection in biological evolution. Based on this entropy increase and the associated increase of statistical probability, we conjecture that the entropy increase is a selection principle for FPCs considered to be variable in earlier discussions.
II Why filling the condensed matter gap is important
@@
Second, the uncertainty and insufficient development of physical models related to higher-energy processes limits our understanding of FPCs Adams (2019). On the other hand, condensed matter physics involves a larger variety of ways to test condensed matter theories. Many of these tests are more readily available than those in high energy physics. As a result, theories and concepts in condensed matter physics have been well verified and cross-checked using a great variety of experiments and simulations.
@@
In the next section, we discuss several examples where FPCs set observability and operation of several key condensed matter effects and phenomena.
III Observability and operation of phase transitions. Emergence of new states with higher entropy
III.1 Fundamental and environmental parameters
We start with separating two types of parameters. The first type are FPCs. The second type are the environmental parameters such as temperature $T$ and pressure $P$ at typical planetary conditions where we do our observations. These parameters are not fundamental but are due to historical accidents Weinberg (1993) such as the distance between the planet and its star which affects the planet’s temperature. In the Solar system, for example, $T$ is in the range 102–103 K.
@@ III.1 Fundamental and environmental parameters
FPCs discussed in this paper will include those central to condensed matter physics: the Planck constant $\hbar$, electron mass $m_{e}$ and charge $e$. These constants define the characteristic scales of distance and energy in condensed matter by setting the Bohr radius $a_{\rm B}$ and Rydberg energy $E_{\rm R}$ as
@@
We will also discuss the mass of the proton (the composite particle), $m_{p}$, which sets the scale of lattice oscillation frequency (see section III.2.2). $m_{p}$ will mostly enter our discussion as the dimensionless ratio
$\beta=\frac{m_{p}}{m_{e}}$ (3)
We note another important dimensionless parameter, the fine-structure constant
$\alpha=\frac{1}{4\pi\epsilon_{0}}\frac{e^{2}}{\hbar c}$ (4)
where $c$ is the speed of light in vacuum. $\alpha$ and $\beta$ play a key role at different length and energy scales in Figure 1, including the stability of particles, nuclear reactions, formation and evolution of stars, synthesis of heavy nuclei and so on. These processes impose constraints on the allowed range of $\alpha$ and $\beta$ (as well as on the strong coupling constant) Barrow (2003); Adams (2019). $\alpha$ will play an indirect role in our discussion in Section III.2.2 because we don’t consider relativistic effects here. This may change if relativistic effects play a role in other condensed matter processes.
This amounts to the following set of FPCs discussed in this paper:
${\rm FPCs}=\{\hbar,e,m_{e},\frac{m_{p}}{m_{e}},\alpha\}$ (5)
III.2 Structural phase transitions
III.2.1 Entropy increase due to a phase transition
Our first case study are structural phase transitions in solids. Let’s consider common silica SiO2 which has several polymorphs at different pressure and temperature, separated by phase transitions. As a model example, let’s consider two silica phases: quartz and high-temperature $\beta$-cristobalite phase Hammonds et al. (1996); Dove (1997).
The solid entropy in the classical high-temperature regime $T\gg\hbar\omega_{\rm D}$, where $\omega_{\rm D}$ is Debye frequency, is Landau and Lifshitz (1970):
@@ III.2.1 Entropy increase due to a phase transition
As expected for high-temperature phases on general grounds Landau and Lifshitz (1970), cristobalite has a more symmetric and open structure compared to quartz. As a result, cristobalite has a significantly larger number of low-frequency phonons, rigid-unit modes, that propagate without the distortion of rigid SiO4 tetrahedra Hammonds et al. (1996); Dove (1997). This gives smaller $\bar{\omega}$ in cristobalite and larger entropy in Eq. (6).
@@
Frequency softening above the phase transition can be seen more generally in terms of the order parameter $\eta$. Considering that phonon frequencies are renormalised by anharmonicity related to a phase transition, the renormalised frequencies $\tilde{\omega}_{i}$ are Dove (1997, 2003):
@@
The entropy increase can also take place below $T_{c}$ if the high-temperature phase is a long-lived metastable state below $T_{c}$. In our quartz-cristobalite example, the phase transition is a reconstructive transition with large energy barriers separating the two phases. As a result, cristobalite is a long-lived metastable phase which can be cooled back to room or any other temperature below $T_{c}$ Hammonds et al. (1996); Dove (1997). We note here that the entropy and other thermodynamic functions of non-equilibrium metastable systems are defined in the commonly used sense: at times shorter than the relaxation time of the non-equilibrium system, this system can be conditionally considered in an equilibrium state where observed properties are described by equilibrium thermodynamics Landau and Lifshitz (1970). This is discussed in Section III.5 in more detail.
@@
The quartz-cristobalite example is readily generalised to many high-temperature polymorphic phases of different solids emerging as a result of phase transitions, with overall entropy increased.
III.2.2 Observability of phase transitions. Entropy changes.
We now observe that if we consider different values of FPCs, the phase transition would not operate. The upper bound on the maximal phonon frequency $\omega_{m}$ (“cage rattling” frequency) can be written in terms of FPCs Trachenko et al. . This bound follows from combining Eq. (2) and the ratio of $\hbar\omega_{m}$ and cohesive energy $E$. The ratio involves a harmonic approximation for the relation between $E$ and $\omega_{m}$ and reads Trachenko and Brazhkin (2020); Trachenko (2023a):
@@
$\omega_{m}$ has the upper bound, $\omega_{m}^{u}$, for $m=m_{p}$ as Trachenko et al.
@@
$\omega_{m}^{u}$ interestingly contains the proton-to-electron mass ratio $\beta$ (3). As mentioned earlier, this ratio is of particular importance for different processes and reactions involving different length scales in Figure 1 Barrow (2003); Adams (2019).
Let’s consider what would happen to a phase transition were FPCs to take different values. Let’s consider variations of FPCs which increase $\omega_{m}$ and $\omega_{m}^{u}$ in Eqs. (13)-(14). Note that this question is well-posed because it is possible to vary $\omega_{m}^{u}$ by varying $m_{e}$ and $m_{p}$ or $\hbar$ and $e$ while keeping $\beta$ and $\alpha$ in Eqs. (3) and (4) unchanged and hence maintain all preceding essential processes (e.g. stellar formation and evolution, nuclear synthesis, stability of ordered structures and so on) intact Barrow (2003); Adams (2019).
@@
Recalling that $\omega_{m}$ is close to $\omega_{\rm D}$, condition (15) implies $T\ll\hbar\omega_{\rm D}$. This, in turn, implies that the high-temperature classical result (6) does not apply because the system is now in the quantum regime. In this regime, instead of Eq. (6), the entropy is Landau and Lifshitz (1970)
@@
The observability of phase transitions can also be discussed on the basis of $T_{c}$. Phase transition theories, including the Ising model and the anharmonic phonon theory (equivalent to Landau theory), expectedly predict that the order of magnitude of $T_{c}$ is set by the energy of interaction between particles Dove (1997). This energy is close to the phonon energy $\hbar\omega_{\rm D}$ or $\hbar\omega_{m}$. Using Eq. (13),
@@
$T_{c}$ can be estimated by recalling that $A^{\frac{1}{2}}$ varies across the periodic table in the range of about 1-15, with an average of 8. Using this value and $\omega_{m}^{u}=3680$ K (14), Eq. (18) gives $T_{c}\approx 500$ K. This is close to $T_{c}$ derived using typical parameters of the anharmonic phonon model and Landau theory Dove (1993) and is of the right order of magnitude of $T_{c}$ typically observed in the range $10^{2}$-$10^{3}$ K.
@@
In summary, the discussion in Section III.2 has so far given us several insights:
@@
    FPCs set the observability and operation of entire physical phenomena such as phase transitions (more examples are in Sections III.3, III.4 and III.5).
@@
    FPCs, by enabling phase transitions, increase the phase space and entropy (more discussion is in Sections III.4, III.5 and IV below).
III.3 Superconducting transition
Perhaps unexpectedly at this point, the results in the previous section lead to the question of observability of superconductivity. We have recently asked Trachenko et al. whether there is a fundamental limit to how high the superconducting temperature, $T_{c}$, can get? Fundamental limit here specifically means the limit in terms of FPCs.
Using the Migdal-Eliashberg theory, the upper limit to $T_{c}$, $T_{c}^{u}$, can be written as $T_{c}^{u}=C\omega_{m}^{u}$, where $\omega_{m}^{u}$ is given in Eq. (14) Trachenko et al. . $C$ is a function of the electron-phonon coupling constant $\lambda$ and initially increases with $\lambda$ but has a maximum around $\lambda=1-2$ because larger $\lambda$ result in instabilities which lower $T_{c}$. This range of $\lambda$ corresponds to $C=0.1-0.2$. Using Eq. (14), this gives $T_{c}^{u}$ as Trachenko et al.